OpenStudy (anonymous):

1/4(3x-16)+1/4x=5 1/3 what is x?

OpenStudy (whovianchick):

1/4(3x-16)+1/4x=5 1/3 3/4x-4+1/4x=5 1/3 x-4= 5 1/3 x=9 1/3